# Surface Reconstruction from Point Clouds via Grid-based Intersection Prediction

### 21 Mar 2024 v1

@@
Abstract
Surface reconstruction from point clouds is a crucial task in the fields of computer vision and computer graphics. SDF-based methods excel at reconstructing smooth meshes with minimal error and artifacts but struggle with representing open surfaces. On the other hand, UDF-based methods can effectively represent open surfaces but often introduce noise near the surface, leading to artifacts in the mesh. In this work, we propose a novel approach that directly predicts the intersection points between sampled line segments of point pairs and implicit surfaces. This method not only preserves the ability to represent open surfaces but also eliminates artifacts in the mesh. Our approach demonstrates state-of-the-art performance on three datasets: ShapeNet, MGN, and ScanNet. The code will be made available upon acceptance.
Keywords:
@@ Abstract
Figure 1: The idea and reconstruction quality of GIFS [28], $\alpha=0.5$, ours and ground-truth, where $\alpha=\frac{|P_{1}Q|}{|P_{1}P_{2}|}$. We can see that ours reconstructs surface with better details and less artifacts.
1 Introduction
@@
Surface reconstruction from point clouds is a long-standing and essential task that has been studied for many years. It plays a crucial role in various modern applications, including visual navigation and robotics. Numerous notable works [13, 15, 14, 7, 17, 1, 2, 27, 23, 24, 20, 25] have significantly advanced this field. Among the traditional methods, Poisson Surface Reconstruction [11], a well-known approach, has yielded good results. However, its performance may deteriorate in the absence of normals and faces with open-surface challenges. With the emergence of deep learning, many researchers have sought to enhance reconstruction results by leveraging the robust representation and predictive capabilities of deep neural networks. These methods can generally be categorized into SDF-based and UDF-based approaches in terms of geometry representation. UDF-based methods excel at representing open surfaces, whereas SDF-based methods struggle in this regard. Yet, predicting UDF values near the surface often introduces noise, despite the critical role these values play in achieving high-quality surface reconstruction. In this study, a novel approach is proposed to enhance surface reconstruction by directly predicting the intersection points between surfaces and line segments. This method not only improves reconstruction quality but also maintains the capacity to represent open surfaces effectively.
@@
Most methods for reconstructing surfaces from point clouds ultimately rely on Marching Cubes or its variants. Marching Cubes require two key pieces of information to generate a triangle mesh. As depicted in Fig. 2 (b), the first piece is the sign of the eight corners in a cube, while the second is the intersection points between the surface and the edges of the cube, such as the exact position of point P after determining that P lies on line AB in Fig. 2 (b). Therefore, in order to achieve better surface reconstruction, it is crucial to enhance the accuracy of both the sign of the eight corners in a cube and the exact intersection positions. While most researchers agree on the importance of the accuracy of the sign of the eight corners in a cube, the significance of the exact intersection position is often overlooked. However, the second aspect is equally important. This can be illustrated through a simple comparison. As demonstrated in Fig. 1, the shape in the fourth column is reconstructed using Marching Cubes with ground-truth signs, and all intersection points are precisely in the middle of the cube edges. On the other hand, the shape in the fifth column is reconstructed with ground-truth signs and ground-truth intersection positions. It is evident that the shape in the fifth column appears smoother, whereas the one in the fourth column exhibits numerous artifacts. These artifacts mainly stem from incorrect intersection positions when $\alpha=0.5$, where $\alpha=\frac{|P_{1}Q|}{|P_{1}P_{2}|}$ in Fig. 1. Therefore, the accuracy of intersection positions significantly impacts the quality of surface reconstruction.
[Uncaptioned image] Figure 2: (a) aims to clarify the intersection module symmetry is necessary. (b) indicate the two parts, cube corner sign and intersection position, we need to predict in reconstructing mesh
[Uncaptioned image] Figure 3: When predicting UDF near the zero-level set, it is easy to generate lots errors due to discontinuity of gradient near the zero-level set.
@@
It is a common phenomenon that predicting the UDF in the vicinity of the surface tends to introduce noise. This is due to the discontinuity of the UDF gradient near the surface. When using a neural network to represent the UDF, accurately predicting the sharp changes near the zero-level set becomes challenging. This issue is illustrated in Fig. 3. The CAP method [31] also highlights this challenge. The presence of UDF noise near the surface can lead to incorrect intersection points between the cube edge and the surface, ultimately resulting in artifacts in the reconstructed mesh. This discrepancy explains why SDF-based methods typically yield smoother results compared to UDF-based approaches.
Although UDF can represent a wider range of surfaces than SDF, it tends to introduce noise near the surface due to the discontinuity of its gradient around the surface. Therefore, we have chosen to avoid predicting UDF and instead focus on directly predicting the intersection points between surfaces and cube edges. The pipeline is illustrated in Fig. 4. This approach allows us to maintain the ability to represent open surfaces while generating smoother surfaces with fewer artifacts compared to UDF-based methods. Furthermore, to ensure surface continuity, we have implemented a special design to predict intersections with a property similar to an odd function. In summary, we can list our main contribution here,
@@
  • Developed a specialized module designed to predict these intersections while maintaining the surface’s continuity.
@@
Figure 4: The pipeline of our method. First, we predict the relative sign of the cube. Then, we predict the intersection between the segment of point pair and implicit surface. Finally, we construct the mesh using the template in Marching Cube.
2 Related Works
@@
Point cloud reconstruction has been a persistent challenge in the fields of graphics and computer vision. Among the traditional methods, Poisson surface reconstruction [11] and ball-pivoting reconstruction [3] stand out as the most significant. The Poisson surface reconstruction method categorizes the query points based on the Poisson indicator function, while the ball-pivoting method generates a continuous surface by simulating a ball rolling over the points. Although these approaches yield satisfactory surface reconstructions, there is still room for further enhancement in performance.
@@
Prominent global methods in this field include DeepSDF[21] and BSP-Net[6], which extract shape features and reconstruct the surface from these features. On the other hand, ConvOccNet[22], SSR-Net[20], DeepMLS[15], and POCO[4] are representative local methods. ConvOccNet[22] converts point cloud features into voxels and enhances features through volume convolution. SSR-Net[20] extracts point features, maps neighborhood point features to octants, and classifies them accordingly. DeepMLS[15] predicts normal and radius for each point and classifies query points based on the moving least-squares equation.
In addition to global and local methods, there are techniques that combine both global and local information, such as Points2Surf[9]. This method aims to regress the absolute SDF value using local information and determine the sign using global information.
Other SDF-based implicit surface reconstruction methods include SAL[1], SALD[2], and On-Surface Prior[18], which aim to transform explicit representations like point clouds and 3D triangle soups into implicit SDF representations. These methods require unique training processes and network parameters for each 3D model. SAL[1] employs MLP to predict shape SDFs with UDF-based metric supervision, while SALD[2] adds a derivative regularization term. On-Surface Prior[18] utilizes a pre-trained UDF-based network to improve SDF predictions.
@@
UDF has been a focal point in the realm of surface representation due to its ability to capture more general surfaces, including open surfaces and partial scans. Various research efforts have delved into leveraging UDF for implicit surface reconstruction. For instance, NDF [7] utilizes UDF to encode surface information and introduces a method for up-sampling point clouds during mesh reconstruction, employing the Ball Pivoting [3] technique. On the other hand, DUDE [27] adopts a volumetric UDF representation encoded within a neural network to describe shapes, leading to promising results in implicit surface reconstruction tasks. Nevertheless, a common challenge across these approaches is the reliance on 3D ground truth labels for training. Furthermore, a key unresolved issue pertains to the direct extraction of iso-surfaces from UDF. In a different approach, UWED [24] leverages Moving Least Squares (MLS) to convert UDF into a dense point cloud, showcasing an alternative strategy for handling UDF representations in surface reconstruction tasks.
3 Methods
3.1 Overview
@@
In this section, we will initially present the complete pipeline of our method. Subsequently, we will delve into the prediction of intersection points and the pair-based sampling technique. Following that, we will discuss the two distinct components: predicting the sign and the intersection point individually. Lastly, we will elucidate the training loss.
3.2 Architecture
Figure 5: The UDF estimation network. Given the input point cloud $\mathcal{P}$ and the upsampled point cloud $\mathcal{P}_{\text{upsample}}$, the network estimates the per-point UDF projection flow in two stages with the guidance of the predicted approaching vector and the point-wise feature, respectively.
@@ 3.2 Architecture
The network architecture is depicted in a straightforward manner, as illustrated in Fig. 5. Initially, the input point cloud with a total of $N$ points is inputted into the encoder to extract features with a dimension of $N*256$. The encoder comprises 4 transformer layers[30, 29], outputting dimensions of $64$, $256$, $256$, and $256$ respectively. Following a similar approach to GIFS [28], we adopt a paired query point sampling method, generating a total of $M$ pairs. Subsequently, an interpolation layer is employed to derive the features of the $2M$ query points. The network then diverges into two distinct paths. In the upper path, the focus is on predicting the relationship between the two points. The feature of the point pair is fed into the sign module to capture the relationship between the points. A multi-layer perceptron (MLP) is utilized to predict whether the two points exhibit the same sign, with a binary cross-entropy loss function applied. In the bottom path, the aim is to predict the position of the intersection point. The feature of the point pair is inputted into the intersection module, followed by the use of an MLP to predict the relative distance from the starting point to the intersection point. A simple regression loss is employed for this prediction. However, in cases where no intersection occurs between the cube edge and the surface, this term is disregarded in the loss calculation.
3.3 Pair Prediction
@@
To sample point pairs for training and testing, we start by considering a point cloud $\mathcal{P}=\{p_{1},p_{2},\cdots,p_{N}\}$. During training, we uniformly sample points around $\mathcal{P}$ at varying scales and ratios. For each point $p_{i}$, a query point $q=p_{i}+U(-1,1)*s$ is randomly sampled, where $s$ takes values of 0.005, 0.01, and 0.02, corresponding to sample ratios of 0.6, 0.3, and 0.1 respectively. Subsequently, a cube with an edge length of 1/256 and $q$ as the upper-left corner is generated, allowing us to sample 12 point pairs corresponding to the 12 edges of the cube. Following this, we determine the intersection point of each edge with the surface and assess whether the point pairs lie on the same side of the surface.
@@
During testing, to reconstruct the entire surface within the given space, we begin by identifying the bounding box of the shape and dividing it into cube meshes with an edge length of 1/256. For each cube within the bounding box, we predict the sign of the cube’s corners and the intersection positions. Utilizing the template matching method in Marching Cubes, we can then generate the triangle mesh. After sampling, we will elaborate on the specific predictions made by our network. Illustrated in Fig. 5, the lower pathway involves selecting a point pair AB, with point $A$ designated as the starting point and $B$ as the endpoint. The intersection point is denoted as $P$, and our aim is to forecast the proportionate distance from $A$ to $P$, expressed as $\frac{|AP|}{|AB|}$. Furthermore, as depicted in Fig. 5, within the upper pathway, we predict whether the point pairs lie on the same or different sides of the surface. The sign module’s name suggests that we solely predict the relationship between the corners of the cube and the upper-left corner within the cube. It is not the sign in the global view but a relative sign to the upper-left corner.
3.4 Sign Module
@@
In the sign module, we formalize the problem as follows: given a start point $A$ and an end point $B$ with features $f_{A}$ and $f_{B}$ respectively, we denote the sign module as $G(f_{A},f_{B})$. The sign module $G(f_{A},f_{B})$ is required to exhibit the following property: when the positions of $A$ and $B$ are exchanged, $G(f_{A},f_{B})$ should vanish, i.e., $G(f_{A},f_{B})=G(f_{B},f_{A})$. This is because the relationship between $A$ and $B$ remains unchanged when their positions are swapped. To maintain this property and enhance the expressive capacity of $G(\cdot,\cdot)$, we design the sign module as follows:
$G(f_{A},f_{B})=concat(f_{A}+f_{B}+pos,f_{A}*f_{B}+pos,max(f_{A},f_{B})+pos),$ (1)
@@
where $pos$ is the positional encoding[26] of $A-B$. Here, in order to keep the property $G(f_{A},f_{B})=G(f_{B},f_{A})$, we abandon the traditional sin-cos positional encoding and adopt only cosine positional encoding. Specifically, $pos(A,B)=cos(\frac{A-B}{2^{i}})$, $i\in\{1,2,\cdots,19\}$.
3.5 Intersection Module

### 09 Apr 2024 v2

@@
Abstract.
Surface reconstruction from point clouds is a crucial task in the fields of computer vision and computer graphics. SDF-based methods excel at reconstructing smooth meshes with minimal error and artefacts but struggle with representing open surfaces. On the other hand, UDF-based methods can effectively represent open surfaces but often introduce noise, leading to artefacts in the mesh. In this work, we propose a novel approach that directly predicts the intersection points between line segment of point pairs and implicit surfaces. To achieve it, we propose two modules named Relative Intersection Module and Sign Module respectively with the feature of point pair as input. To preserve the continuity of the surface, we also integrate symmetry into the two modules, which means the position of predicted intersection will not change even if the input order of the point pair changes. This method not only preserves the ability to represent open surfaces but also eliminates most artefacts on the mesh. Our approach demonstrates state-of-the-art performance on three datasets: ShapeNet, MGN, and ScanNet. The code will be made available upon acceptance.
@@ Abstract.
journal: TOGjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Computing methodologies Reconstructionccs: Computing methodologies Shape representations
1. Introduction
@@
Surface reconstruction from point clouds is a long-standing and essential task that has been studied for many years. It plays a crucial role in various modern applications, including visual navigation and robotics. Numerous notable works (Li et al., 2022; Liu et al., 2021; Liao et al., 2018; Chibane et al., 2020b; Ma et al., 2020; Atzmon and Lipman, 2020a, b; Venkatesh et al., 2020; Wang et al., 2022a; Richa et al., [n. d.]; Zhenxing et al., 2020; Zhao et al., 2023) have significantly advanced this field. Among the traditional methods, Poisson Surface Reconstruction (Kazhdan et al., 2013), a well-known approach, has yielded good results. However, its performance may deteriorate in the absence of normals and faces with open-surface challenges. With the emergence of deep learning, many researchers have sought to enhance reconstruction results by leveraging the robust representation and predictive capabilities of deep neural networks. These methods can generally be categorized into SDF-based (Signed Distance Function) and UDF-based (Unsigned Distance Function) approaches in terms of geometry representation. UDF-based methods excel at representing open surfaces, whereas SDF-based methods struggle in this regard. Yet, predicting UDF values often introduces noise, despite the critical role these UDF values play in achieving high-quality surface reconstruction. In this study, a novel approach is proposed to enhance surface reconstruction by directly predicting the intersection points between surfaces and line segments. This method not only improves reconstruction quality but also maintains the capacity to represent open surfaces effectively.
@@
In particular, the motivation of our method comes from the following two aspects. First, Most methods for reconstructing surfaces from point clouds ultimately rely on Marching Cubes or its variants. Marching Cubes requires two key pieces of information to generate a triangle mesh. As depicted in Fig. 3 (a), the first piece is the sign of the vertices at the eight corners of the cube, while the second is the intersection points between the surface and the edges of the cube, as the exact position of point P after determining that P lies on line AB in Fig. 3 (b). Therefore, in order to achieve better surface reconstruction, it is crucial to enhance the accuracy of both the sign of the vertices at the eight corners of a cube and the exact intersection positions. While most researchers agree on the importance of the accuracy of the sign of the vertices at the eight corners of a cube, the significance of the exact intersection position is often overlooked. However, the second aspect is also very important. This can be illustrated through a simple comparison. As demonstrated in Fig. 2, the shape in the left column is reconstructed using Marching Cubes with ground-truth signs, and all intersection points are randomly chosen in $(0,1)$, supposing the length of the cube edge is $1$. On the other hand, the shape in the right column is reconstructed with ground-truth signs and ground-truth intersection positions. It is evident that the shape in the right column appears smoother, whereas the one in the left column exhibits numerous artefacts. These artefacts mainly stem from incorrect intersection positions when $\alpha$ is disturbed by the random noise, where $\alpha=\frac{|AP|}{|AB|}$ in Fig. 1. Therefore, the accuracy of intersection positions significantly impacts the quality of surface reconstruction.
Figure 2. The left column is shape visualization when we keep the vertices sign at the corner of cube correct and give the intersection a random disturbing. The right column is the GT shape.
@@
Secondly, it is a common phenomenon that predicting the UDF in the vicinity of the surface tends to introduce noise. This noise arises from the discontinuity of the UDF gradient on the two side of the surface. When utilizing a neural network to represent the UDF, accurately predicting the sharp changes in the UDF gradient around the zero-level set becomes a challenging task. Furthermore, we provide a small yet compelling example to illustrate this point in Section 4.1. The CAP method (Zhou et al., 2022) also emphasizes this challenge. The presence of UDF noise can result in inaccurate intersection points between the cube edge and the surface, leading to artefacts in the reconstructed mesh. This discrepancy elucidates why SDF-based methods typically produce smoother results compared to UDF-based approaches. Building upon these insights, we aim to propose a method that can not only reconstruct the open surface but also mitigate the impact of UDF noise.
Therefore, we have chosen to avoid predicting UDF and instead focus on directly predicting the intersection points between surfaces and cube edges. Specifically, we have designed two modules: the Relative Sign Module and the Intersection Module.
In the Relative Sign module, we predict whether two points have the same sign and incorporate a symmetrical design in the module. The symmetry guarantees that the relative sign will not change when the input order of point pair changes. Once we confirm that the signs of the two points are different, we proceed to use the Intersection Module.
In the Intersection Module, we directly predict the intersection points between the cube edges and the surface, while also implementing a symmetrical design. The symmetry aims to achieve the following purpose. In the point pair with A as start point and B as end point, the intersection point will not change when A and B exchange their position. The symmetry of both the Intersection Module and the Relative Sign module ensures the continuity of the surface.
To demonstrate the superiority of our approach, we conducted evaluations using chamfer distance and normal consistency metrics across three diverse datasets: ShapeNet, MGN, and ScanNet, encompassing watertight shapes, open surfaces, and partially scanned scenes. Compared to state-of-the-art methods, our method consistently outperforms them, yielding the best results across all three datasets. Notably, our method excels on ShapeNet, demonstrating a remarkable 30% reduction of $CD_{1}$ and a 1.8% enhancement of $NC$.
In summary, we can list our main innovative points,
@@
  • Developed two specialized modules designed to predict these intersections while maintaining the surface’s continuity.
@@
2. Related Works
@@
Point cloud reconstruction has been a persistent challenge in the fields of graphics and computer vision. Among the traditional methods, Poisson surface reconstruction (Kazhdan et al., 2013) and ball-pivoting reconstruction (Bernardini et al., 1999) stand out as the most significant. The Poisson surface reconstruction method categorizes the query points based on the Poisson indicator function, while the ball-pivoting method generates a continuous surface by simulating a ball rolling over the points. Although these approaches yield satisfactory surface reconstructions, there is still room for further enhancement in performance.
@@
Early representative works include AtlasNet(Groueix et al., 2018), IF-Net(Chibane et al., 2020a), DeepSDF(Park et al., 2019) and BSP-Net(Chen et al., 2020), which extract shape features and reconstruct the surface from these features. Later, ConvOccNet(Peng et al., 2020), SSR-Net(Zhenxing et al., 2020), DeepMLS(Liu et al., 2021; Kolluri, 2008), and POCO(Boulch and Marlet, 2022), O-CNN(Wang et al., 2020) DeepCurrents(Palmer et al., [n. d.]), LP-DIF(Wang et al., 2023a) and Venkatesh(Venkatesh et al., 2021) develop this methods. ConvOccNet(Peng et al., 2020) converts point cloud features into voxels and enhances features through volume convolution. SSR-Net(Zhenxing et al., 2020) extracts point features, maps neighborhood point features to octants, and classifies them accordingly. DeepMLS(Liu et al., 2021; Kolluri, 2008) predicts normal and radius for each point and classifies query points based on the moving least-squares equation. ALTO(Wang et al., 2023b; Hu et al., 2020) dynamically adjusts the representation of local geometry. Dual Octree(Wang et al., 2022a) proposes a new graph convolution operator defined over a regular grid of features fused from irregular neighboring octree nodes at different levels, simultaneously improving the performance and reducing the cost. Li, T(Li et al., 2022) propose a dynamic code to represent the local geometry with dynamic resource adaptively. Points2Surf(Erler et al., 2020) aims to regress the absolute SDF value using local information and determine the sign using global information.
Other SDF-based implicit surface reconstruction methods include SAL(Atzmon and Lipman, 2020a), SALD(Atzmon and Lipman, 2020b), StEiK(Yang et al., 2023), SIREN(Sitzmann et al., 2020), DiGS(Ben-Shabat et al., [n. d.]) and On-Surface Prior(Ma et al., [n. d.]), which aim to transform explicit representations like point clouds and 3D triangle soups into implicit SDF representations. These methods require unique training processes and network parameters for each 3D model. SAL(Atzmon and Lipman, 2020a) employs MLP to predict shape SDFs with UDF-based metric supervision, while SALD(Atzmon and Lipman, 2020b) adds a derivative regularization term. On-Surface Prior(Ma et al., [n. d.]) utilizes a pre-trained UDF-based network to improve SDF predictions.
Differential Marching Cube methods is related to ours, including Deep Marching Cube(Liao et al., 2018) and Neural Marching Cube(Chen and Zhang, 2021). Both of them are based on SDF.
@@
UDF has been a focal point in the realm of surface representation due to its ability to capture more general surfaces, including open surfaces and partial scans. Various research efforts have delved into leveraging UDF for implicit surface reconstruction. For instance, NDF (Chibane et al., 2020b) utilizes UDF to encode surface information and introduces a method for up-sampling point clouds during mesh reconstruction, employing the Ball Pivoting (Bernardini et al., 1999) technique. On the other hand, DUDE (Venkatesh et al., 2020) adopts a volumetric UDF representation encoded within a neural network to describe shapes, leading to promising results in implicit surface reconstruction tasks. Nevertheless, a common challenge across these approaches is the reliance on 3D ground truth labels for training. Furthermore, a key unresolved issue pertains to the direct extraction of iso-surfaces from UDF. In a different approach, UWED (Richa et al., [n. d.]) leverages Moving Least Squares (MLS) to convert UDF into a dense point cloud, showcasing an alternative strategy for handling UDF representations in surface reconstruction tasks. MeshUDF(Guillard et al., 2021)vote the sign of local cube corner according the local geometry. HSDF(Wang et al., 2022b) adopts the hybrid SDF and UDF to represent open surface.
Figure 3. This figure indicates the two parts in reconstructing mesh, the sign of vertices of 8 corners in a cube in (a) and the intersection position in (b).
3. Methods
3.1. Overview
Figure 4. The pipeline of our method. First, we predict the relative sign of the vertices at the corners of the cube. Then, we predict the intersection between the line segment of the point pair and the implicit surface. Finally, we construct the mesh using the template in Marching Cubes.
@@ 3.1. Overview
In this section, we will initially present the complete pipeline of our method. Subsequently, we will delve into the prediction of intersection points and the pair-based sampling technique. Following that, we will discuss the two distinct components: predicting the relative sign and the intersection point individually. Lastly, we will elucidate the training loss.
3.2. Architecture
@@
The total pipeline is shown in Fig. 4. The network architecture is depicted in a straightforward manner, as illustrated in Fig. 5. Initially, the input point cloud with a total of $N$ points is inputted into the encoder to extract features with a dimension of $N*256$. The encoder comprises 4 transformer layers(Zhao et al., 2021, 2020), outputting dimensions of $64$, $256$, $256$, and $256$ respectively. Following a similar approach to GIFS (Ye et al., 2022), we adopt a paired query point sampling method, generating a total of $M$ pairs. Subsequently, an interpolation layer is employed to derive the features of the $2M$ query points. The network then diverges into two distinct paths. In the upper path, the focus is on predicting the relative sign of the two points. The feature of the point pair is fed into the Relative Sign Module to capture the relationship between the points. A multi-layer perceptron (MLP) is utilized to predict whether the two points exhibit the same sign, with a binary cross-entropy loss function applied. In the bottom path, the aim is to predict the position of the intersection point. The feature of the point pair is inputted into the Intersection Module, followed by the use of a MLP to predict the relative distance from the starting point to the intersection point. A simple regression loss is employed for this prediction. However, in cases where no intersection occurs between the cube edge and the surface, this term is disregarded in the loss calculation.
3.3. Pair Prediction
@@
To sample point pairs for training and testing, we start by considering a point cloud $\mathcal{P}=\{p_{1},p_{2},\cdots,p_{N}\}$. During training, we uniformly sample points around $\mathcal{P}$ at varying scales and ratios. For each point $p_{i}$, a query point $q=p_{i}+U(-1,1)\times s$ is randomly sampled, where $s$ takes values of 0.005, 0.01, and 0.02, corresponding to sample ratios of 0.6, 0.3, and 0.1 respectively. Subsequently, a cube with an edge length of 1/256 and $q$ as the upper-left-front corner is generated, allowing us to sample 12 point pairs corresponding to the 12 edges of the cube. Following this, we determine the intersection point of each edge with the surface and assess whether the point pairs lie on the same side of the surface.
@@
During testing, to reconstruct the entire surface within the given space, we begin by identifying the bounding box of the shape and dividing it into cube meshes with an edge length of 1/256. For each cube within the bounding box, we predict the relative sign of the cube’s corners and the intersection positions, where the reference point is the vertex at the upper-left-front corner of the cube. Utilizing the template matching method in Marching Cubes, we can then generate the triangle mesh. After sampling, we will elaborate on the specific predictions made by our network. Illustrated in Fig. 5, the lower pathway involves selecting a point pair AB, with point $A$ designated as the starting point and $B$ as the end point. The intersection point is denoted as $P$, and our aim is to forecast the proportionate distance from $A$ to $P$, expressed as $\frac{|AP|}{|AB|}$. Furthermore, as depicted in Fig. 5, within the upper pathway, we predict whether the point pairs lie on the same or different sides of the surface. The Relative Sign Module’s name suggests that we solely predict the relationship between the corners of the cube and the upper-left-front corner within the cube. It is not the sign in the global view but a relative sign to the upper-left-front corner.
3.4. Relative Sign Module
@@
In the Relative Sign Module, we formalize the problem as follows: given a start point $A$ and an end point $B$ with features $f_{A}$ and $f_{B}$ respectively, we denote the Relative Sign Module as $G(f_{A},f_{B})$. The Relative Sign Module $G(f_{A},f_{B})$ is required to exhibit the following property: when the positions of $A$ and $B$ are exchanged, $G(f_{A},f_{B})$ should vanish, i.e., $G(f_{A},f_{B})=G(f_{B},f_{A})$. This is because the relationship between $A$ and $B$ remains unchanged when their positions are swapped. To maintain this property and enhance the expressive capacity of $G(\cdot,\cdot)$, we design the Relative Sign Module as follows:
(1) $G(f_{A},f_{B})=concat(f_{A}+f_{B}+pos,f_{A}*f_{B}+pos,max(f_{A},f_{B})+pos),$
@@
where $pos$ is the positional encoding(Tancik et al., 2020) of $A-B$. Here, in order to keep the property $G(f_{A},f_{B})=G(f_{B},f_{A})$, we abandon the traditional sin-cos positional encoding and adopt only cosine positional encoding. Specifically, $pos(A,B)=cos(\frac{A-B}{2^{i}})$, $i\in\{0,2,\cdots,19\}$.
Figure 5. The network architecture of our method, where $M$ is the number of point pair, $N$ is the number of points in a point cloud.
3.5. Intersection Module